The sum of three numbers is 85.
The second is 7 more than the first.
The third is 2 more than four times the second. Find the numbers

Answer by Edwin McCravy(20060)

The sum of three numbers is 85.
(1) x+y+z = 85

The second is 7 more than the first.
(2) y = x+7

The third is 2 more than four times the second.
(3) z = 4y + 2

Find the numbers
Plug (2) into (3)

(3) z = 4y+2
    z = 4(x+7)+2
    z = 4x+28+2
(4) z = 4x+30

Plug (2) and (4) into (1)

(1) x+y+z = 85
    x+(x+7)+(4x+30) = 85
    x+x+7+4x+30 = 85
    6x+37 = 85
    6x = 48
(5)  x = 8 

Plug (5) in (2)

(2) y = x+7
    y = 8+7
(6) y = 15

Plug (6) in (3)

(3) z = 4y + 2
    z = 4(15) + 2
    z = 60 + 2
    z = 62

Edwin